40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 8 20:11:54 2012, Jenne, Update, Alignment, AS steering moved out of POY's way, 2 green beams onto PSL table AS_8Nov2012.pngREFL_8Nov2012.png
    Reply  Thu Nov 8 20:54:08 2012, Manasa, Ayaka, Update, Alignment, Reconfirming on IPPOS, IPANG and oplevs centering 
    Reply  Fri Nov 9 08:11:57 2012, Steve, Update, Alignment, nothing moved - all good opsingGOOD.pngop8h.png8amF.pngkeepfailing.png
       Reply  Fri Nov 9 11:38:38 2012, jamie, Update, General, we're closing up 
          Reply  Fri Nov 9 17:15:05 2012, Manasa, Steve, Ayaka, Update, General, We're closed! Pumping down monday morning 
Message ID: 7692     Entry time: Fri Nov 9 08:11:57 2012     In reply to: 7689     Reply to this: 7693
Author: Steve 
Type: Update 
Category: Alignment 
Subject: nothing moved - all good 

 Oplevs and IP ANG are still centered. Why  do the SRM and PRM move 5X  more ?  I could not check the sensing voltages because the computer failed.

AS and REFL are looking the same as last night.

ALL LOOKING GOOD!

 

Attachment 1: opsingGOOD.png  33 kB  | Hide | Hide all
opsingGOOD.png
Attachment 2: op8h.png  59 kB  | Hide | Hide all
op8h.png
Attachment 3: 8amF.png  18 kB  | Hide | Hide all
8amF.png
Attachment 4: keepfailing.png  25 kB  | Hide | Hide all
keepfailing.png
ELOG V3.1.3-